OpenStudy (anonymous):

(2x + 5y)(2x − 5y) solve please

OpenStudy (anonymous):

(a-b)(a+b)=a^2-b^2 so (2x + 5y)(2x − 5y)=(2x)^2-(5y)^2=4x^2-25y^2

OpenStudy (jackellyn):

(2x)(2x)+(-5y)(2x)+(5y)(2x)+(5y)(-5y) 4x^2 -10xy + 10xy - 25y^2 4x^2-25y^2

OpenStudy (anonymous):

4x^2-25y^2